OpenStudy (anonymous):

Use substitution to solve for Y in the system of equations. 2x-2y=20 4x+y=20 A.-4 B.-6 C.0 D.4

OpenStudy (anonymous):

I will give medals!!!

OpenStudy (kidrah69):

isolate the equation 2 since it has a y inly new equation would be y=-4x+20 now sub it ibot equation 1 and solve. 2x-2(-4x+20)=20

OpenStudy (kidrah69):

*into

OpenStudy (anonymous):

Thank you!

OpenStudy (kidrah69):

oo we are not done yet!! ;-; this will only find our x value :) I want you to solve the equation I gave you find the x value

OpenStudy (kidrah69):

2x+8x-40=20 combine like terms. 10x=20 10/10=0 20/10/? once you find the x value plug it into equation 2 :) y=-4(x value goes here)+20 y=? there you go thats all there is :) welcome to openstudy!

OpenStudy (anonymous):

Thank you! But I already did haha. I am not stupid lol